OpenStudy (anonymous):

What is the sum of a 56–term arithmetic sequence where the first term is 6 and the last term is 391? 10,719 11,116 11,513 11,910

OpenStudy (raden):

just use the formula : --------------- Sn = n/2*(a+an) --------------- S56 = 56/2*(6+391) S56 = 28*(397) S56 = 11,116